Consolidated Pasta is currently expected to pay annual dividends of $10 a share in perpetuity on the 1.2 million shares that are outstanding. Shareholders require a 8% rate of return from Consolidated stock.

a. What is the price of Consolidated stock? (Do not round intermediate calculations.)

b. What is the total market value of its equity? (Enter your answer in millions.) Consolidated now decides to increase next year’s dividend to $20 a share, without changing its investment or borrowing plans. Thereafter the company will revert to its policy of distributing $10 million a year.

c. How much new equity capital will the company need to raise to finance the extra dividend payment? (Enter your answer in millions.)

d. What will be the total present value of dividends paid each year on the new shares that the company will need to issue? (Enter your answer in millions.)

e. What will be the transfer of value from the old shareholders to the new shareholders? (Enter your answer in millions.)
Answer #1

a. Stock price = Dividend per share/ Rate of return = $10/0.08 = $125

b. Market value of equity = Stock price * No of shares = $125 * 1.2 million shares = $150 million

c. New equity = Extra dividend * No of shares = $10 * 1.2 million shares = $12 million

d. The present value of the new shares will need to be $12 million, since that is the amount that of the investment.

e. Transfer value from the old shareholders to the new shareholders is exactly $12 million. This is the amount of dividend received and reflects the new money raised.